OpenStudy (anonymous):

Find the sum of all integers 30 to 80, INCLUSIVE. (Is there a formula to do this?)

OpenStudy (anonymous):

((30+80)*50)/2

OpenStudy (anonymous):

it will be 30, 31, 32,... 80 forms A.P. Sn=n/2[2a+(n-1)d] i.e. 50/2[2*30+(50-1)1] i.e. 25[60+49] i.e.25*109 =2725

OpenStudy (anonymous):

@myko and @priyadabas answers are my answer chocies, but which one is right?

OpenStudy (anonymous):

bouth

OpenStudy (anonymous):

Nope yours equal 2750

OpenStudy (anonymous):

ohh, so then it's mine, :)

OpenStudy (anonymous):

-_- not really it could be wrong

OpenStudy (unklerhaukus):

you could find halfway between 30 and 80, which is 55 and add 30+80, 31+79,....54+56, 55 that is 25 times 110 +55

OpenStudy (unklerhaukus):

=2805

OpenStudy (anonymous):

30 31 32 33 34..... 80 80 79 78 77 76 .... 30 ----------------------- + = twice the summ

OpenStudy (anonymous):

OMG The answer chcies are: A. 2725, what @priyadabas got B. 2750, what @myko got C. 2805, what @UnkleRhaukus got D. 2835

OpenStudy (anonymous):

SO CONFUSED :(

OpenStudy (anonymous):

30 to 80, INCLUSIVE

OpenStudy (anonymous):

The formula for calculating sum of an A.P. is Sn=n/2[2a+(n-1)d] where n=no. of terms a=first term d=diff. b/w two terms.

OpenStudy (anonymous):

BUT WHO IS RIGHT??? I dont even know the answer...

OpenStudy (unklerhaukus):

myko forgot to add the 55, the middle term

OpenStudy (anonymous):

Alright, so it leaves down to @priyadabas and @UnkleRhaukus

OpenStudy (anonymous):

SO CONFUSED because the answer chocies are: A. 2725, what @priyadabas got B. 2750, what @myko got C. 2805, what @UnkleRhaukus got D. 2835

OpenStudy (unklerhaukus):

priyadabas has left off the last term, 80

OpenStudy (anonymous):

So... @UnkleRhaukus is right?

OpenStudy (anonymous):

@UnkleRhaukus - U're wrong. Last term is called tn. I'm right.

OpenStudy (anonymous):

AH!!! SO CONFUSED!!!

OpenStudy (anonymous):

n it's of no use in the sum only difference n no. of terms are used.

OpenStudy (unklerhaukus):

@priyadabas can you see an error in my method?

OpenStudy (anonymous):

@UnkleRhaukus -U've used the ncert formula- n(n+1)/2 wich is not wrong but it can't be used like that here. It is for the sum of 'n' integers.

OpenStudy (anonymous):

((30+80)*51)/2=2805 i made mistake befor, becouse there are 51 terms to sum, not 50

OpenStudy (anonymous):

@Spyush

OpenStudy (unklerhaukus):

but \[(30+80)\times50)/2+55=2805\] \[(30+80)\times51)/2+55=2860\]

OpenStudy (anonymous):

55=110/2

OpenStudy (unklerhaukus):

now im confused

OpenStudy (anonymous):

@UnkleRhaukus's result 2805 including both left and right ends [ 30, 80 ]

OpenStudy (anonymous):

It is 2805. You first find the average \[( 30+80)\div2=55\] Then find the number of numbers so \[80-30+1=51\] Now you multiply these two numbers together \[55 \times 510=2805\]

OpenStudy (anonymous):

@myko what?

OpenStudy (kinggeorge):

I'm getting 2805. \[\sum_{n=30}^{80}n=\sum_{n=1}^{80}n-\sum_{n=1}^{29}n\]\[\frac{80(81)}{2}-\frac{29(30)}{2}=40(81)-29(15)=3240-435=2805\]

OpenStudy (anonymous):

now there is a method i can understand!

OpenStudy (anonymous):

and an answer that is correct

OpenStudy (callisto):

I love the formula \(1+2+3+...+ n = \frac{n}{2}(1+n)\) Just saying..